OpenStudy (sparklestaraa):

The two lines, P and Q, are graphed below: (graph below) Determine the solution and the reasoning that justifies the solution to the systems of equations. (4 points) (0, 8), because this point is true for both the equations (0, 8), because the lines intersect the y-axis at these points (-2, 4), because this point makes both the equations incorrect (-2, 4), because it is the point of intersection of the two graphs

OpenStudy (sparklestaraa):

I think its D @butterflydreamer

OpenStudy (butterflydreamer):

correct! :D

OpenStudy (sparklestaraa):

A system of equations is shown below: 6x - 2y = 3 (equation 1) 5x + 3y = 4 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof? Show that the solution to the system of equations 10x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 10x - 2y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x - y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations I think this one is B @butterflydreamer

OpenStudy (butterflydreamer):

6x + 5x = 11x true? So this means that options A and B are knocked out. So now we have decide between option C and D :) So what is 6x + 5y -2y +3y = 3 + 4

OpenStudy (sparklestaraa):

ohhhhhhh D :D because that has plus :D

OpenStudy (butterflydreamer):

yepp :) What they wanted you to do was add equation 1 and equation 2 because the question already mentions the bit about the "sum" and since equation 2 is multiplied by 1, it will remain unchanged :) so D is correct

OpenStudy (sparklestaraa):

ok I'm going to tag u in a few more :D

OpenStudy (butterflydreamer):

xD i would love to stay but i really need to head off :( I have an early day for school tmrw. Unless your questions are really quick !

OpenStudy (sparklestaraa):

They will be really quick because i just need to check the answers

OpenStudy (butterflydreamer):

okay! fire away.